OpenStudy (anonymous):

Choose the correct simplification of x to the 2nd power over y to the 5th power all raised to the 4th power.

OpenStudy (zzr0ck3r):

\((\dfrac{x^2}{y^5})^4\) hinte: \((\dfrac{a}{b})^c=\dfrac{a^c}{b^c}\) and \((a^b)^c=a^{bc}\)

OpenStudy (zzr0ck3r):

hint*